# Quantifying Misalignment Between Agents: Towards a Sociotechnical Understanding of Alignment

### v2

@@ 1 Introduction
In recent years, growing concerns have emerged about the AI alignment problem (e.g. Russell 2019; Christian 2020). With respect to the current AI systems that exist today, Russell (Russell 2019) made the bold yet compelling claim that social media AI today is already misaligned with humanity (e.g. through extensive disinformation spread). As agentic AI systems are growing in importance and increasingly deployed (Kapoor et al. 2024), concerns about misalignment of agent-based systems have grown (Chan et al. 2023; Gabriel et al. 2024). Yet with few exceptions (e.g. Critch and Krueger 2020), much prior work focuses on single AI settings (Ji et al. 2024), on value specification and learning (Critch and Krueger 2020; Leike et al. 2018), or centers technosolutionist fixes. A healthier alternative would train a sociotechnical lens (Lazar and Nelson 2023) on multi-agent settings (Gabriel et al. 2024), which may include a complex mix of AI agents and humans, potentially holding a dizzying array of competing and often conflicting goals. Critch and Krueger posed the question, “where could one draw the threshold between ‘not very well aligned’ and ‘misaligned’[..]?” (2020, pg. 14), while Lazar and Nelson recently called for a “sociotechnical approach to AI safety,” stating unequivocally that “no group of experts (especially not technologists alone) should unilaterally decide what risks count, what harms matter, and to which values safe AI should be aligned” (2023). In order to gain a deeper understanding of alignment of AI in its sociotechnical environment, we need to expand our definitions of alignment towards a sociotechnical, agent-based understanding of misalignment that accounts for the complexity of real-world systems in an increasingly AI-agentic world. The question emerges: with whom are agents aligned or misaligned, and on what areas? The alignment of an AI agent or system might be with an adversary, rather than with the developers and/or owners of it; and the alignment—or lack thereof—might be context-, user-, or agent-dependent.
@@
    We introduce a novel mathematical model of misalignment which accounts for a population of human and AI agents, and which is parameterized by problem areas, allowing any pair of agents to be simultaneously both aligned and misaligned. Our model affords explanatory power for the complexity of real-world, sociotechnical settings in which AI agents are deployed.
@@ 2 Related Work
The deployment of powerful agentic AI systems has led to increased concerns about misalignment (Chan et al. 2023; Gabriel et al. 2024). Several researchers argue that existing AI systems already exacerbate threats to information ecosystems and collective decision-making (Bucknall and Dori-Hacohen 2022; Russell 2019; Bak-Coleman et al. 2021; Seger et al. 2020). However, most prior work defines misalignment as a global characteristic of an AI system, often as a binary (Sierra et al. 2021). Recent work in machine ethics attempts to answer questions like “whom should AI align with?”, while others measure alignment with respect to “human values” (Ji et al. 2024). However, while some researchers have noted alignment issues posed by value pluralism (Gabriel 2020), methods for measuring (mis)alignment between different cultures and AI agents remain underdeveloped.
In order to expand our understanding on misalignment to address this gap, our work draws significant inspiration from a computational model of contention in human populations proposed by Jang, Dori-Hacohen, and Allan (2017). Their model quantifies the proportion of people in disagreement on stances regarding a topic, parameterized by the observed group of individuals. We extend this approach in several key ways: (1) we adapt the model to capture misalignment with respect to goals, rather than stances on topics; (2) we extend the population to include both human and AI agents, allowing for analysis of mixed-agent scenarios; (3) we introduce the concept of “problem areas” to segment and analyze alignment across different domains of interaction; (4) our model allows for varying levels of goal conflict and importance weights, providing a more nuanced representation of alignment dynamics. While the contention paper focused on controversy in public discourse (Jang, Dori-Hacohen, and Allan 2017), our model provides a framework for quantifying misalignment in complex sociotechnical systems where humans and AI agents interact. This approach offers greater explanatory power for real-world scenarios and potential applications in AI alignment and safety research.
@@ 3 Modeling misalignment in populations
Our approach to modeling misalignment rests on a key observation: understanding and ”solving” the AI alignment problem requires first grappling with the challenges of aligning humans. Human conflicts are ubiquitous, highlighting the difficulty of alignment even without AI agents. Building on the computational model for contention (Jang, Dori-Hacohen, and Allan 2017), we adapt the approach to capture misalignment with respect to goals. Our model allows for a hybrid population of human and AI agents, quantifying misalignment based on individuals’ goals in problem areas.
Definitions and Notation. We define $\mathcal{A}=\{a_{1}..a_{n}\}$ as a set of $n$ agents and $\mathcal{P}=\{p_{1}..p_{m}\}$ as a set of $m$ problem areas of interest to at least one agent in $\mathcal{A}$. $h(a,g,j)$ is a relationship denoting that agent $a$ holds goal $g$ in problem area $j$. $g^{i}_{j}$ is the goal held by agent $i$ in problem area $j$, and $w^{i}_{j}$ is the weight, representing the importance that agent $i$ assigns to problem area $j$.
@@
$h(a,g_{0},j)\iff\not\exists g\in\hat{\mathcal{G}_{j}}\text{ s.t. }h(a,g,j).$
@@
Measuring Conflict. We introduce a measure, $c$, to denote the incompatibility of a pair of goals:
  • $c(g_{x},g_{y})=1$: Goals $g_{x}$ and $g_{y}$ are in complete conflict; they are mutually exclusive.
  • $c(g_{x},g_{y})=0$: Goals $g_{x}$ and $g_{y}$ are completely compatible and aligned.
By construction, $c(g_{x},g_{x})=0$ and $c(g_{x},g_{0})=0$.
Goal Groups. We define a goal group as a subset of agents that holds the same goal:
$\mathcal{A}_{g}=\{a\in\mathcal{A}|h(a,g,j)\}$
@@
Quantifying Misalignment. Assuming agents regard each $p$ with equal weight, we can quantify misalignment as the probability that two randomly selected agents will hold incompatible goals:
$Pr(0|\mathcal{A},p):=$ $\,\,Pr(a_{1},a_{2}\text{ selected randomly from }\mathcal{A})$
$\cdot c(g^{1}_{p},g^{2}_{p})$
The multiplication of the probability and conflict components calculates an expected conflict between two randomly selected agents’ goals. By doing this for all possible pairs of agents, we obtain an overall measure of misalignment in the population for a given problem area. Note that the equation assumes that all agents are equally likely to be selected; and that the conflict function $c(\cdot,\cdot)$ allows for varying degrees of conflict between goals, rather than just binary conflict/no-conflict situations. Finally, the equation provides a single scalar value representing the overall misalignment in the population for a specific problem area. The benefit of this approach is that it affords a simple representation capturing the real-world phenomena of “strange bedfellows,” where a pair agents may be highly aligned in one specific area despite being highly misaligned in others.
Mutually Exclusive Goals. To simplify our analysis, we can consider scenarios with mutually exclusive goals. This approach allows for a more straightforward quantification of misalignment. We introduce the following constraints:
@@
    Every agent holds only one goal in a problem area.
@@
    Each goal is in conflict with each other goal.
@@
    A lack of a goal ($g_{0}$) is not in conflict with any explicit goal.
Under these constraints, we can compute $Pr(0|\mathcal{A},p)$ as follows:
$Pr(0|\mathcal{A},p)=\frac{\sum_{g\in\mathcal{G}}\sum_{g^{\prime}\in\mathcal{G},g^{\prime}<g}2|\mathcal{A}_{g}||\mathcal{A}_{g^{\prime}}|}{|\mathcal{A}|^{2}}$
Where $|\mathcal{A}_{g}|$ is the number of agents holding goal $g$ and $|\mathcal{A}|$ is the total number of agents. The alignment probability is then simply $Pr(1|\mathcal{A},p)=1-Pr(0|\mathcal{A},p)$.
This equation allows us to derive a parametric quantity for misalignment of uniformly distributed agents in the range $[0,\frac{|\hat{\mathcal{G}}|-1}{|\hat{\mathcal{G}}|}]$.
@@
Overall Misalignment Across Problem Areas. To determine the misalignment of $\mathcal{A}$ across all problem areas, we take the arithmetic mean of $Pr(0|\mathcal{A},p)$ across all $p$’s for that $\mathcal{A}$. Some problem areas are more important to us than others, and this importance changes depending on the agent. To aggregate misalignment accordingly, let each agent have a weight parameter $w$ for each problem area, such that 0 is the minimum amount an agent can care about a problem area, and 1 is the maximum amount. Then, multiply each problem area’s misalignment by the geometric mean of the sampled agents’ weights for that $p$. This gives us the following equation:
$Pr(0|\mathcal{A},\mathcal{P}):=$
$\frac{1}{|\mathcal{P}|}\sum_{p\in\mathcal{P}}$ $\Big{(}Pr(a_{1},a_{2}\text{ selected randomly from }\mathcal{A})$
$\cdot c(g^{1}_{p},g^{2}_{p})\cdot\sqrt{w^{1}_{p}\cdot w^{2}_{p}}\Big{)}$
@@
Interpreting Misalignment Scores. The misalignment score provides insights into the degree of goal conflict within a set of agents:
@@ 5 Results
Effect of Population Size and Problem Areas
Figure 1 illustrates how misalignment evolves as we increase the number of agents, with varying numbers of problem areas. The results reveal a consistent pattern across different numbers of problem areas. Initially, there is high variance in misalignment due to the random distribution of goals among a small number of agents. However, as the population size increases, the misalignment converges to a stable value. Notably, this convergence point is independent of the number of problem areas, because the overall misalignment is calculated as the arithmetic mean across all areas and each problem area has the same population misalignment. This finding demonstrates that our model captures the intuitive notion that larger populations tend to stabilize in their overall misalignment, even when individual problem areas may have different early dynamics. It suggests that in complex multi-agent systems, the aggregate level of misalignment becomes more predictable as the number of agents grows.
Impact of Goal Count
Figure 2 shows how the number of goals in each problem area affects misalignment. The experiment reveals a clear relationship between the number of goals and the level of misalignment. As the number of goals per problem area increases, so does the overall misalignment. For a problem area with with $\hat{\mathcal{G}}$ non-zero goals, the misalignment approaches $(\hat{\mathcal{G}}-1)/\hat{\mathcal{G}}$ as the number of agents grows. This pattern aligns with the intuitive understanding that a greater number of mutually exclusive goals leads to higher potential for misalignment within a population. The result highlights our model’s ability to capture the complexity of multi-goal scenarios, reflecting the increased potential for conflict as the number of distinct goals grows. It suggests that in real-world situations, systems with a higher number of competing objectives are likely to exhibit greater levels of misalignment, emphasizing the importance of goal prioritization and conflict resolution in multi-agent environments.
Figure 1: Varying Problem Areas. Random goal assignment, 3 goals per problem area, 100 runs per data point, max weight and conflict values. Null goals disallowed.
@@
Weight Sensitivity
Figure 3 examines the effect of changing weights for a single goal group in one problem area, while keeping other weights at maximum. When the weight of one goal group is reduced to zero in a single problem area, it effectively eliminates misalignment for that specific area. This demonstrates the model’s ability to capture scenarios where certain objectives become irrelevant or unimportant to a subset of agents. As more problem areas are added to the simulation, the impact of weight changes in a single area diminishes, reflecting the model’s capability to balance multiple concerns. Interestingly, when the number of goals per problem area is increased to four, the misalignment starts at a higher level and approaches 0.75, consistent with the $(\hat{\mathcal{G}}-1)/\hat{\mathcal{G}}$ pattern observed in the previous experiment. These findings showcase our model’s sensitivity to goal weights and its ability to capture nuanced interactions between problem areas. They suggest that in complex systems, the relative importance assigned to different goals can significantly influence overall alignment, and that this effect is modulated by the number and nature of the problem areas under consideration.
@@
Goal Distribution Effects
Figure 4 shows how misalignment changes as we vary the proportion of agents assigned to different goals. The experiment reveals dynamics of misalignment in populations with varying goal allocations. Notably, misalignment reaches its peak when goals are evenly distributed among agents. Specifically, for a problem area with $\hat{\mathcal{G}}$ non-zero goals, the maximum misalignment occurs when the proportion of agents assigned to any single goal is $1/\hat{\mathcal{G}}$. As the population becomes more homogeneous in its goals - that is, as a larger proportion of agents adopt the same goal - the overall misalignment decreases. This pattern holds true regardless of the total number of goals, though the peak misalignment value increases with the number of available goals. The results suggest (in line with our intuitions) that in real-world scenarios, populations with a diverse and balanced set of objectives may experience higher levels of misalignment compared to those with a more dominant goal or consensus.
Conflict Levels and Problem Area Interactions
@@
The analysis of conflict levels and problem area interactions reveals complex dynamics in our misalignment model. As expected, misalignment increases with both the number of goals and the level of conflict between goals. However, the effect of multiple problem areas is more nuanced, depending on the average conflict across those areas rather than their individual values. Notably, scenarios with the same average conflict level converge to similar misalignment values, regardless of how that average is achieved across different problem areas. This finding suggests that our model can effectively capture the aggregate effect of conflicts across multiple domains, providing a holistic view of misalignment in complex environments. More broadly, these results demonstrate our model’s capability to handle intricate scenarios with varying conflict levels across multiple problem areas, offering a sophisticated tool for analyzing misalignment in multi-faceted, real-world situations where conflicts may be unevenly distributed across different domains of interaction.
Thus, our simulation experiments validate key properties of our misalignment model, demonstrating its ability to capture intuitive aspects of multi-agent alignment while revealing non-obvious dynamics in complex scenarios. With these findings as a solid foundation, we now move on to applying our model to realistic case studies.

### v3

@@ 1 Introduction
In recent years, growing concerns have emerged about the AI alignment problem (e.g. Russell 2019; Christian 2020). With respect to the current AI systems that exist today, Russell (Russell 2019) made the bold yet compelling claim that social media AI today is already misaligned with humanity (e.g. through extensive disinformation spread). As agentic AI systems are growing in importance and increasingly deployed (Kapoor et al. 2024), concerns about misalignment of agent-based systems have grown (Chan et al. 2023; Gabriel et al. 2024). Yet with few exceptions (e.g. Critch and Krueger 2020), much prior work focuses on single AI settings (Ji et al. 2024), on value specification and learning (Critch and Krueger 2020; Leike et al. 2018), or centers technosolutionist fixes. A healthier alternative would train a sociotechnical lens (Lazar and Nelson 2023) on multi-agent settings (Gabriel et al. 2024), which may include a complex mix of AI agents and humans, potentially holding a dizzying array of competing and often conflicting goals. Critch and Krueger posed the question, “where could one draw the threshold between ‘not very well aligned’ and ‘misaligned’[..]?” (2020, pg. 14), while Lazar and Nelson call for a “sociotechnical approach to AI safety,” stating unequivocally that “no group of experts (especially not technologists alone) should unilaterally decide what risks count, what harms matter, and to which values safe AI should be aligned” (2023). In order to consider the alignment of AI in its sociotechnical environment, we need to expand our definitions of alignment toward a sociotechnical, agent-based understanding that accounts for the complexity of real-world systems in an increasingly AI-agentic world. The question emerges: with whom are agents aligned or misaligned, and on what areas? The alignment of an AI agent or system might be with an adversary, rather than with the developers and/or owners of it; and the alignment—or lack thereof—might be context-, user-, or agent-dependent.
@@
    We introduce a novel mathematical model of misalignment which accounts for a population of human and AI agents, and which is parameterized by problem areas, allowing any pair of agents to be simultaneously aligned and misaligned. Our model affords explanatory power for the complexity of real-world, sociotechnical settings in which AI agents are deployed.
@@ 2 Related Work
The deployment of powerful agentic AI systems has led to increased concerns about misalignment (Chan et al. 2023; Gabriel et al. 2024). Several researchers argue that existing AI systems already exacerbate threats to information ecosystems and collective decision-making (Bucknall and Dori-Hacohen 2022; Dori-Hacohen et al. 2021; Russell 2019; Bak-Coleman et al. 2021; Seger et al. 2020). However, most prior work defines misalignment as a global characteristic of an AI system, often as a binary (Sierra et al. 2021). Recent work in machine ethics attempts to answer questions like “whom should AI align with?”, while others measure alignment with respect to “human values” (Ji et al. 2024). However, while some researchers have noted alignment issues posed by value pluralism (Gabriel 2020), methods for measuring (mis)alignment between different cultures and AI agents remain underdeveloped.
To expand our understanding of misalignment to address this gap, our work draws significant inspiration from a computational model of contention in human populations proposed by Jang, Dori-Hacohen, and Allan (2017). Their model quantifies the proportion of people in disagreement on stances regarding a topic, parameterized by the observed group of individuals. We extend this approach in several key ways: (1) we adapt the model to capture misalignment with respect to goals, rather than stances on topics; (2) we extend the population to include both human and AI agents, allowing for analysis of mixed-agent scenarios; (3) we introduce the concept of “problem areas” to segment and analyze alignment across different domains of interaction; (4) we allow for varying levels of goal conflict and importance, providing a nuanced representation of alignment dynamics. While the contention paper focused on controversy in public discourse (Jang, Dori-Hacohen, and Allan 2017), our model provides a framework for quantifying misalignment in complex sociotechnical systems where humans and AI agents interact. This approach offers greater real-world explanatory power, with applications in both AI research and policy (see §C).
@@ 3 Modeling misalignment in populations
Our approach to modeling misalignment rests on a key observation: understanding and ”solving” the AI alignment problem requires first grappling with the challenges of aligning humans. The ubiquity of human conflicts highlights the difficulty of alignment even without AI. Adapting the computational contention model’s individualized framing (Jang, Dori-Hacohen, and Allan 2017), we quantify misalignment based on each human and AI agent’s goals in problem areas.
3.1  Definitions and Notation. We define $\mathcal{A}=\{a_{1}..a_{n}\}$ as a set of $n$ agents and $\mathcal{P}=\{p_{1}..p_{m}\}$ as a set of $m$ problem areas of interest to at least one agent in $\mathcal{A}$. $h(a,g,j)$ is a relationship denoting that agent $a$ holds goal $g$ in problem area $j$. $g^{i}_{j}$ is the goal held by agent $i$ in problem area $j$, and $w^{i}_{j}$ is the weight, representing the importance that agent $i$ assigns to problem area $j$.
@@
$h(a,g_{0},j)\iff\not\exists g\in\hat{\mathcal{G}_{j}}\text{ s.t. }h(a,g,j).$ (1)
@@
3.2  Measuring Conflict. In order to capture the notion that goals can be compatible or conflicting, we introduce $c$, the probability that a pair of goals is incompatible:
  • $Pr(c=1|g^{1}_{p},g^{2}_{p})=1$: Goals $g^{1}_{p}$ and $g^{2}_{p}$ are in complete conflict; they are mutually exclusive.
  • $Pr(c=1|g^{1}_{p},g^{2}_{p})=0$: Goals $g^{1}_{p}$ and $g^{2}_{p}$ are completely compatible and aligned.
For readability, we use $c(g^{1}_{p},g^{2}_{p})$ as shorthand for $Pr(c=1|g^{1}_{p},g^{2}_{p})$.Since $c$ represents the probability of conflict between goals, it is a real number bounded in the range $[0,1]$. By construction, $c(g^{x}_{p},g^{x}_{p})=0$ and $c(g^{x}_{p},g^{0}_{p})=0$.
3.3  Goal Groups. We define a goal group as a subset of agents that holds the same goal:
$\mathcal{A}_{g}=\{a\in\mathcal{A}|h(a,g,j)\}$ (2)
@@
3.4  Quantifying Misalignment. In a single problem area, we can quantify misalignment as the probability that two randomly selected agents will hold incompatible goals:
$Pr(1|\mathcal{A},p):=$ $\,\,Pr(a_{1},a_{2}\text{ selected randomly from }\mathcal{A},$ (3)
$a_{1}\neq a_{2})\cdot c(g^{1}_{p},g^{2}_{p})$
The multiplication of the sampling and conflict components calculates the probability that any two randomly selected agents’ goals will be in conflict. By doing this for all possible pairs of agents, we obtain an overall measure of misalignment in the population for a given problem area. The equation assumes the following constraints:
@@
    Each agent holds no more than one goal per problem area.
@@
    A lack of a goal ($g_{0}$) is not in conflict with any explicit goal.
@@
    All agents are equally likely to be selected.
We discuss the practical limitations of these constraints and how to remove them in Appendix A. Note that the conflict function $c(\cdot,\cdot)$ allows for varying degrees of conflict between goals, rather than just binary conflict/no-conflict situations. Finally, the equation provides a single real scalar in the range $[0,1]$ representing the overall misalignment in the population for a specific problem area. The benefit of this approach is that it affords a simple representation capturing the real-world phenomena of “strange bedfellows,” where a pair of agents may be highly aligned in one specific area despite being highly misaligned in others.
3.5  Mutually Exclusive Goals. To simplify our analysis, we can consider scenarios with mutually exclusive goals, meaning that each goal in $p$ completely conflicts with each other goal in $p$. This constraint allows for a more straightforward quantification of misalignment as follows:
$Pr(1|\mathcal{A},p)=\frac{\sum_{g\in\mathcal{G}}\sum_{g^{\prime}\in\mathcal{G},g^{\prime}<g}2|\mathcal{A}_{g}||\mathcal{A}_{g^{\prime}}|}{|\mathcal{A}|(|\mathcal{A}|-1)}$ (4)
Where $|\mathcal{A}_{g}|$ is the number of agents holding goal $g$ and $|\mathcal{A}|$ is the total number of agents. The alignment probability is then simply $Pr(0|\mathcal{A},p)=1-Pr(1|\mathcal{A},p)$.
This equation allows us to derive a parametric quantity for misalignment of uniformly distributed agents in the range $[0,\frac{|\hat{\mathcal{G}}|-1}{|\hat{\mathcal{G}}|})$. See §5.1, §5.2, and §5.4 for experimental confirmation of this bound, and Appendices B.3 and B.4 for mathematical proof and interpretation.
@@
3.6  Overall Misalignment Across Problem Areas. To find the misalignment of $\mathcal{A}$ across all problem areas, we take the arithmetic mean of $Pr(1|\mathcal{A},p)$ across all $p$’s for that $\mathcal{A}$. Some problem areas are more important to us than others, and this importance changes depending on the agent. To aggregate misalignment accordingly, let each agent have a weight parameter $w$ for each problem area, such that 0 is the minimum amount an agent can care about a problem area, and 1 is the maximum amount. Then, multiply each problem area’s misalignment by the geometric mean of the sampled agents’ weights for that $p$. This gives us the following equation:
$Pr(1|\mathcal{A},\mathcal{P}):=$ (5)
$\frac{1}{|\mathcal{P}|}\sum_{p\in\mathcal{P}}$ $\Big{(}Pr(a_{1},a_{2}\text{ selected randomly from }\mathcal{A}$
$a_{1}\neq a_{2})\cdot c(g^{1}_{p},g^{2}_{p})\cdot\sqrt{w^{1}_{p}\cdot w^{2}_{p}}\Big{)}$
@@
3.7  Interpreting Misalignment Scores. The misalignment score provides insights into the degree of goal conflict within a set of agents:
@@ 5 Results
5.1  Varying Problem Areas.
Figure 1 illustrates how misalignment evolves as we increase the number of agents, with varying numbers of problem areas. The results reveal a consistent pattern across different numbers of problem areas. Initially, there is high variance in misalignment due to the random distribution of goals among a small number of agents. However, as the population size increases, the misalignment converges to a stable value. Notably, this convergence point is independent of the number of problem areas, because the overall misalignment is calculated as the arithmetic mean across all areas and each problem area has the same population misalignment. This finding demonstrates that our model captures the intuitive notion that larger populations tend to stabilize in their overall misalignment, even when they may have different early dynamics. It suggests that in complex multi-agent systems, the aggregate level of misalignment becomes more predictable as the number of agents grows and their goal distribution settles.
5.2  Varying Goals.
Figure 2 shows that as the number of conflicting goals per problem area increases, so does the overall misalignment. For a problem area with with $\hat{\mathcal{G}}$ non-zero goals, the misalignment approaches $\frac{\hat{\mathcal{G}}-1}{\hat{\mathcal{G}}}$ as the number of agents grows. This pattern aligns with the intuitive understanding that a greater number of mutually exclusive goals leads to higher potential for misalignment within a population. The result highlights our model’s ability to capture the complexity of multi-goal scenarios, reflecting the increased potential for conflict as the number of distinct goals grows. It suggests that in real-world situations, systems with a higher number of competing objectives are likely to exhibit greater levels of misalignment, showing the importance of goal prioritization and conflict resolution in multi-agent environments.
Figure 1: Varying Problem Areas. Random goal assignment, 3 goals per problem area, 100 runs per data point, max weights mutually exclusive and conflict values. Null goals disallowed.
@@
5.3  Weight Sensitivity.
Figure 3 examines the effect of changing weights for a single goal group in one problem area, while keeping other weights at maximum. When the weight of one goal group is reduced to zero in a single problem area, it effectively eliminates misalignment for that specific area. This aspect of the model captures scenarios where some objectives are irrelevant or unimportant to a subset of agents. As more problem areas are added to the simulation, the impact of weight changes in a single area diminishes, reflecting the model’s capability to balance multiple concerns. Interestingly, when the number of goals per problem area is increased to four, the misalignment starts at a higher level and approaches 0.75, consistent with the $\frac{\hat{\mathcal{G}}-1}{\hat{\mathcal{G}}}$ pattern observed in the previous experiment. These findings show that our model is sensitive to goal weights and can capture nuanced interactions between problem areas. They suggest that in complex systems, the relative importance assigned to different goals can significantly influence overall alignment, and that this effect is modulated by the number of problem areas considered.
@@
5.4  Goal Distribution.
Figure 4 shows how the dynamics of misalignment change as we vary the proportion of agents assigned to different goals. Notably, misalignment reaches its peak when goals are evenly distributed among agents. Specifically, for a problem area with $\hat{\mathcal{G}}$ non-zero goals, the maximum misalignment occurs when the proportion of agents assigned to any single goal is $\frac{1}{\hat{\mathcal{G}}}$. As the population becomes more homogeneous in its goals - that is, as a larger proportion of agents adopt the same goal - the overall misalignment decreases. As the Goal 1 group shrinks, agents are redistributed, so the leftmost value of each curve is the peak value of the curve with one less goal. This pattern holds regardless of the total number of goals, though the peak misalignment value increases with the number of available goals. The results suggest (intuitively) that in real-world scenarios, divided populations with few common objectives are more misaligned than to those with a more dominant consensus.
5.5  Conflict Levels.
@@
As expected, misalignment increases with both the number of goals and the level of conflict between goals. However, the effect of multiple problem areas is more nuanced, depending on the average conflict across those areas rather than their individual values. Notably, scenarios with the same average conflict level converge to similar misalignment values, regardless of how that average is achieved across different problem areas. This finding suggests that our model effectively captures the aggregate effect of conflicts across multiple domains, providing a holistic view of misalignment. More broadly, this demonstrates our model’s ability to handle intricate scenarios with varying conflict levels across multiple problem areas, offering a sophisticated tool for analyzing misalignment in multi-faceted, real-world situations where conflicts may be unevenly distributed across different domains of interaction.
Together, our experiments validate key properties of our misalignment model, demonstrating its ability to capture intuitive aspects of multi-agent alignment while revealing nontrivial dynamics in complex scenarios. With this solid foundation, we now apply our model to realistic case studies.